How do you find experimetal probability?
Experimental probability is calculated by taking the data produced from a performed experiment and calculating probability from that data. An example would be flipping a coin. The theoretical probability of landing on heads is 50%, .5 or 1/2, as is the theoretical probability of landing on tails. If during an experiment, however, a coin is flipped 100 times and lands on heads 60 times and tails 40 times, the experimental probability for this experiment for landing on heads is 60%, .6 or 6/10. The experimental probability of landing on tails would be 40%, .4, or 6/10.
it depends on the total number of marbles you have!
Is it possible for the standard deviation of a set of data to be zero?
Yes, but only if every element in the data set is exactly the same. Therefore, very unlikely.
What is the theoretical probability of flipping a coin 60 times?
Theoretical is 50% Heads, 50% tails: 30-Heads, 30-Tails (theoretical)
How many possible outcomes of tossing a coin and rolling a 6 sided number cube are there?
Coin - 1/2 chance for either heads or tails
Die - 16 chance of landing on 1, 2, 3, 4, 5, or 6
The worth of word 13 cards of McDonalds Limited Edition 1993 Game Day collector cards can vary from person to person. There are people out there who may be looking for these. The best way to find out how much they are worth is to do a quick eBay look, and see if anyone is selling theirs, and how much. You can also find a private collector or appraiser to help you find the worth. It is hard to find the worth of a product without visually looking at it.
The probability of getting two tails in the first two is 1/4. And it does not matter how many more times the coins are tossed after the first two tosses.
Why is 7 the most likely answer If you roll 2 fair dice and add the scores?
If you roll two dice, the following reuslts are possible:
2: 1+1
3: 1+2 , 2+1
4: 1+3, 2+2, 3+1
5: 1+4, 2+3, 3+2, 4+1
6: 1+5, 2+4, 3+3, 4+2, 5+1
7: 1+6, 2+5, 3+4, 4+3, 5+2, 6+1
8: 2+6, 3+5, 4+4, 5+3, 6+2
9: 3+6, 4+5, 5+4, 6+3
10: 4+6, 5+5, 6+4
11: 5+6, 6+5
12: 6+6
As you can see, the greatest number of permutations result in a total of 7. Its probability is 6/36 or 1/6.
What is the probability of rolling a pair of dice and obtaining a sum of 9?
Assuming of course that we are talking about cubic dice.
combinations that make 9:
6&3, 5&4, 4&5, 3&6. There are 4 combinations that make 9.
There are 36 possible combination of two dice.
So 4/36 is 1/9
What is the simplest form of 3 out of 8?
3 out of 8 is already in simplest form, because 3 is prime and 3 and 8 are not divisible by each other. In decimal form, 3 out of 8 is 0.375.
How many different ways can the letters in the word annually be arranged?
There are 8!/(2!*2!*2!) = 5040 ways.
What is the number of distinguishable permutations of the letters in the word EFFECTIVE?
The number of permutations of the letters EFFECTIVE is 9 factorial or 362,880. Since the letter E is repeated twice we need to divide that by 4, to get 90,720. Since the letter F is repeated once we need to divide that by 2, to get 45,360.
What is the probability of getting an even number on a dice?
3/6 or 50%. if simplified it would equal 1/2.
What is the probability of getting three heads in a row if you flip a coin five times?
Each flip can land in 2 ways. Total number of possible ways that 5 flips can land is
2 x 2 x 2 x 2 x 2 = 32 different ways.
There are only 5 different ways to get exactly 3 heads in a row in 5 flips.
They are:
H H H T T
H H H T H
T H H H T
T T H H H
H T H H H.
The probabilty is 5/32 = 15.625 percent.
====================================================
That's if you mean exactly 3 heads in a row ... no less and no more.
What if you asked for "at least 3 heads in a row" ?
There are 8 ways to do that:
The original five ways to get 3 in a row . . .
H H H T T
H H H T H
T H H H T
T T H H H
H T H H H
Plus 2 ways to get 4 in a row . . .
T H H H H
H H H H T
Plus a way to get 5 in a row . . .
H H H H H.
The probability of one of these outcomes is 8/32 = 25 percent.
What is the probability of getting heads more than once if you flip a coin three times?
For 3 coin flips:
87% chance of getting heads at least once
25% chance of getting heads twice
13% chance of getting heads all three times
How do you calculate the density of a six sided die?
A die is basically a cube. Calculate density by measuring the volume of the cube and dividing that into the weight of the cube.
The volume can be calculated using the formula for a cube. Let's assume the holes in the faces are negligible. All three sides of a cube(Length, width and height) are the same length.
Volume = length of side X length of side X length of side
Measure the weight of the die.
Density = Weight / Volume
If the dots on the die are cut out, then you can get a little more accurate by measuring the size of the dot and calculating the depth. Then you would try to calculate the volume of a dot on the die.
Then subtract this volume for each of the dots on the faces of the die. There are 21 dots(I think).
Is the different between an individual score and mean of the group of scores is called deviation?
No, it is called the absolute deviation.
The probability of rolling snake eyes on any one throw is 1 in 36. The probability of NOT rolling a snake eyes is 35 in 36. The only outcome here which we do not desire is that in which a non-snake eye roll occurs four times in a row, therefore, we take (35/36)^4 to find the probability of NEVER getting a snake eyes, and then subtract that answer from 1 to get the chances for any outcome WITH at least one pair of snake eyes.
The chances are 10.66% of getting at least one snake eyes in four rolls.
When two dice are rolled find the probability of getting a sum of 5 or 6?
There are 36 permutations of two dice. Of these, 9 have a sum of 5 or 6, so the probability of rolling a sum of 5 or 6 on two dice is 9 in 36, or 1 in 9, or about 0.1111.
How do you get if you roll a die 100 times how many times would you expect to get 1?
If you roll a die 100 times, you would expect to get a 1 about 17 times, because the probability of getting a 1 is 1 in 6, or 0.1667. However, that is theoretical probability; experimental probability - the actual results of doing this 100 times - might not be 17, but if you did this a large number of times, the experimental results would indeed begin to approach the theoretical results.
What is the probability if rolling a sum of 7 or 8 given the second die is a 1 or a 2?
1/4 = 0.25
If the second die is 1 or 2, then there are 12 possible outcomes.
To score 7: 5+2 or 6+1 ⇒ 2 rolls
To score 8: 6+2 ⇒ 1 roll
⇒ to score 7 or 8: 2 + 1 = 3 rolls
So there are 3 rolls to score the required sums out of 12 possible rolls, giving a probability of:
3/12 = 1/4 = 0.25